# Towards Efficient Real-Time Video Motion Transfer via Generative Time Series Modeling

### v1

@@ 2 Related Works
In addition to keypoint-based approaches, recent advancements in video motion transfer have also included proposals based on diffusion-based models. MotionShop [70] introduces Mixture of Score Guidance (MSG), which decomposes motion into separate components and operates directly on pretrained diffusion models without additional training. Similarly, MotionFlow [33] utilizes cross-attention maps to facilitate motion transfer during inference, enabling more flexible and generalizable motion synthesis. Furthermore, DiTFlow [39] optimizes Attention Motion Flow (AMF) to transfer motion by utilizing Diffusion Transformers (DiTs) for higher motion consistency. Although these diffusion-based techniques achieve high-fidelity motion synthesis, their dependence on iterative denoising and complex optimization processes results in significant computational overhead, making them less viable for real-time applications. Compared to diffusion-based approaches, keypoint based methods like FOMM are more computationally efficient, making them ideal for real-time video applications where low-latency motion transfer is crucial. Our study builds upon keypoint based motion transfer methods such as the FOMM and demonstrates high generalizability, thereby enabling efficient bandwidth optimization across various keypoint-based architectures such as [50, 51, 62, 72]. This makes our proposed architecture particularly well-suited for applications such as video reconstruction and animation while ensuring real-time processing capabilities.
@@ 3.1 The Proposed Pipeline
Our real time motion transfer proposal is based on the FOMM pipeline [51] , although in general our method can be used for realizing bandwidth savings in other keypoint based architectures such as [72, 62]. In the FOMM architecture, inputs to the pipeline consist of a source image $\textbf{S}\in\mathbb{R}^{3\times H\times W}$, where 3 denotes the number of channels corresponding to Red, Green, Blue (RGB), and $H$ and $W$ denote the height and width of the image respectively. and a sequence of $N$ driving video frames $\textbf{D}\in\mathbb{R}^{3\times H\times W}$. At the start of the pipeline there is a keypoint detector which is a convolutional neural network that is used to produce $K$ feature maps from the source image S and each frame of the driving video D. These extracted feature maps are then normalized and condensed to form keypoints by computing the spatial expectation of the maps [34]. Each keypoint consists of the coordinates $x,y$ along with parameters of the local affine transformations around each keypoint. The latter set of features model the motion around a keypoint as a $2$x2222 Jacobian matrix. In this work we set the hyperparameter $K=10$ which gives sufficiently good results. The 20 components representing the coordinates and the 40 components of the Jacobian matrix form a 60 dimensional time series. The keypoint predictor takes $N$ such keypoints corresponding to the driving video frames $\bf D$ and forecasts the next $M$ values. Following this the predicted keypoints go through the dense motion network which is used to align the keypoints captured from source image S and driving video D. The dense motion network is another convolutional neural network that generates heatmaps from the keypoints and then computes the dense motion field. The motion field is basically a function $\hat{\mathcal{T}}_{S\leftarrow D}$ which assigns every pixel position in D to its equivalent position in S. In addition, an occlusion mask $\hat{\mathcal{O}}_{S\leftarrow D}$ is generated by the dense motion network that is responsible for determining which image parts of D must be reconstituted by warping from S and which sections should be infilled based on the context. Finally, the generation module synthesizes the source image S moving in accordance with the driving video D. To achieve this, we utilize a generator network that warps the source image based on $\hat{\mathcal{T}}_{S\leftarrow D}$ and inpaints the occluded regions that are not visible in the source image [51]. In this paper we demonstrate the use of the real-time motion transfer pipeline for multi-step ahead single video frame prediction as well as generating diverse samples of future video frames. The proposed pipeline for single and diverse sample prediction are demonstrated in Figures 1 and 2.
@@ 3.2 Keypoint Prediction Models
GRU was proposed in [6] and is a variant of the Recurrent Neural Network (RNN) which is designed to perform forecasting of high dimensional data. The problem of keypoint prediction can also be addressed using Deep Generative Models such as Variational Autoencoder (VAE) [22] and Normalizing Flow (NF) [10]. Both of these are generative models that learn a probabilistic representation of high-dimensional data. Although VAEs and NFs are not inherently architected to perform forecasting, in this work we adapt them for keypoint prediction by reconstructing a future value from a given input. Our primary focus in this work is to demonstrate the forecasting capabilities of two generative time series models namely VRNN and GRU-NF. The VRNN model proposed by [7] combines the strengths of VAEs and RNNs to model sequential data whereas the GRU-NF proposed by [42] combines a GRU with an NF. In the following subsections we review the mathematical setup for each of these five models and also discuss how they are trained in our real-time motion transfer pipeline for keypoint forecasting.
@@ 3.2.1 Gated Recurrent Unit (GRU)
Let an input sequence of keypoints be denoted as $\bm{x}=(\bm{x}_{1},\bm{x}_{2},...,\bm{x}_{t})$. At each timestep $t$, the GRU maintains a hidden state $\bm{h}_{t}$ that is updated based on the current input $\bm{x}_{t}$ and the previous hidden state $\bm{h}_{t-1}$. This update mechanism of the hidden states in GRUs incorporates gating units—specifically, an update gate $\bm{z}_{t}$ and a reset gate $\bm{r}_{t}$—which control the flow of information through the network [6]. First, the reset gate is computed as:
$\bm{r}_{t}=\sigma\left(W_{r}\bm{x}_{t}+U_{r}\bm{h}_{t-1}\right)$ (1)
where $\sigma(\cdot)$ is the sigmoid function, and $W_{r}$, $U_{r}$ are learnable weight matrices. Using the reset gate, a candidate hidden state $\tilde{\bm{h}}_{t}$ is then calculated as below:
$\tilde{\bm{h}}_{t}=\tanh\left(W\bm{x}_{t}+U\left(\bm{r}_{t}\odot\bm{h}_{t-1}\right)\right)$ (2)
@@
$\bm{z}_{t}=\sigma\left(W_{z}\bm{x}_{t}+U_{z}\bm{h}_{t-1}\right)$ (3)
where $W_{z}$, $U_{z}$ are learnable weight matrices. Finally, the hidden state is estimated using the update gate $\bm{z}_{t}$ and the candidate hidden state $\bm{\tilde{h}}_{t}$ as follows:
$\bm{h}_{t}=\left(1-\bm{z}_{t}\right)\odot\bm{h}_{t-1}+\bm{z}_{t}\odot\tilde{\bm{h}}_{t}$ (4)
This gating mechanism allows the GRU to selectively retain or update information from previous input. During prediction, the GRU maps its hidden states to a probability distribution $p$ of outputs $\bm{x}_{t+1}$ as below:
$p\left(\bm{x}_{t+1}\mid\bm{x}_{\leq t}\right)=g_{\tau}\left(\bm{h}_{t}\right)$ (5)
where $g_{\tau}$ is a nonlinear function with learnable parameters $\tau$ that transforms the hidden state $\bm{h}_{t}$ to the corresponding output $\bm{x}_{t+1}$. In our work, the GRU is trained to predict $\bm{x}_{t+1}$ from $\bm{x}_{t}$ by minimizing the mean squared error (MSE) loss for $N$ training samples between the predicted output $\bm{\hat{x}}$ and its corresponding ground truth $\bm{x}$ at time $t+1$ as follows:
$\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}(\hat{\bm{x}}_{t+1}^{(i)}-\bm{x}_{t+1}^{(i)})^{2}$ (6)
@@ 3.2.2 Variational Autoencoder (VAE)
The VAE is a deep generative model which is trained to minimize the reconstruction error between its inputs and outputs while simultaneously learning a structured, continuous latent space that enables smooth interpolation and sampling of new data. A VAE consists of three main components: an encoder that maps the input data to a distribution over latent variables, a latent space that captures uncertainty through probabilistic representations, and a decoder that reconstructs data using samples drawn from this latent distribution. For a time series of keypoints, a VAE encodes the input $\bm{x}_{t}$ at each timestep $t$ into a latent representation $\bm{z}_{t}$ through a probabilistic encoder $q_{\phi}$. The encoder estimates the posterior distribution of the latent variable $\bm{z}_{t}$ as below:
$q_{\phi}(\bm{z}_{t}|\bm{x}_{t})=\mathcal{N}\left(\mu_{\phi}(\bm{x}_{t}),\sigma_{\phi}^{2}(\bm{x}_{t})\right)$ (7)
where $\mu_{\phi}(\bm{x}_{t})$ and $\sigma_{\phi}^{2}(\bm{x}_{t})$ are the learned mean and variance of the Gaussian distribution. To enable backpropagation during training, the latent variable $\bm{z}_{t}$ is sampled using the reparameterization trick [46]:
$\bm{z}_{t}=\mu_{\phi}(\bm{x}_{t})+\sigma_{\phi}(\bm{x}_{t})\odot\epsilon,\quad\epsilon\sim\mathcal{N}(0,I)$ (8)
where $\epsilon$ is a noise vector drawn from a standard normal distribution and $\odot$ denotes element-wise multiplication. The decoder $p_{\theta}$, parameterized by weights $\theta$, reconstructs the input $\bm{x}_{t}$ from the latent variable $\bm{z}_{t}$ by estimating the likelihood as below:
$p_{\theta}(\bm{x}_{t}|\bm{z}_{t})=\mathcal{N}\left(\mu_{\theta}(\bm{z}_{t}),\sigma_{\theta}^{2}(\bm{z}_{t})\right)$ (9)
Here $p_{\theta}(\bm{x}_{t}|\bm{z}_{t})$ represents the likelihood of generating outputs ${\bf x}_{t}$ given the latent variable ${\bf z}_{t}$, the decoder’s output is given by $\mu_{\theta}(\bm{z}_{t})$ and $\sigma_{\theta}^{2}(\bm{z}_{t})$ is the variance associated with the reconstruction. Figure 3 illustrates the architecture of a VAE.
@@
In our work, given $k>1$ as the prediction horizon, the VAE is trained to predict $\bm{x}_{t+k}$ from $\bm{x}_{t}$ at time $t$ using the loss function as given below:
$\mathcal{L}_{\text{VAE}}=\mathcal{L}_{\text{MSE}}+KL[q_{\phi}(\bm{z}_{t}|\bm{x}_{t})||p(\bm{z}_{t})]$ (10)
@@
$\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}(\hat{\bm{x}}_{t+k}^{(i)}-\bm{x}_{t+k}^{(i)})^{2}$ (11)
$KL(q_{\phi}(\bf{z}_{t}|\bf{x}_{t})\,\|\,p(\bf{z}_{t}))$ $=-\frac{1}{2N}\sum_{i=1}^{N}\sum_{j=1}^{J}\Big{[}1+\log\left(\sigma_{\phi,j}^{2}(\bm{x}_{t,i})\right)$ (12)
$\quad\quad-\mu_{\phi,j}^{2}(\bm{x}_{t,i})-\sigma_{\phi,j}^{2}(\bm{x}_{t,i})\Big{]}$
The first term in the loss function minimizes the reconstruction loss over $N$ training samples between the predicted output $\bm{\hat{x}}$ and its corresponding ground truth $\bm{x}$ at time $t+k$ while the second term, $KL[\cdot||\cdot]$, is the Kullback-Leibler divergence between the approximate posterior $q_{\phi}(\bm{z}_{t}\mid\bm{x}_{t})$ and the prior $p(\bm{z}_{t})$ where $\mathcal{N}(0,I)$ is typically chosen as the prior distribution over the latent space.
During inference, the VAE takes as input a sequence of values from timesteps $t_{1},\dots,t_{n}$, and generates the next $k$ future values by reconstructing each $\bm{x}_{k+i}$ from $\bm{x}_{i}$, for $i=1,\dots,n$, where $k>=n$, thereby shifting the input index forward with each prediction.

### v2

@@ 2 Related Works
In addition to keypoint-based approaches, alternative approaches to motion representation include dense optical flow fields [lu2023transflow], prototype-based motion learners [lu2024promotion], and part-based explicit pose modeling [han2024prototypical]. These models provide rich motion encoding, especially in structured environments such as human motion. However, they often rely on supervision, require expensive pretraining, and are less adaptable to non-human motion or real-time applications. Moreover, recent works like [liu2020video] have shown that over-complex motion features may not generalize well in real-time driving applications, reinforcing our choice of lightweight representations. Recent advancements in video motion transfer have also included proposals based on diffusion-based models. MotionShop [Yesiltepe2024] introduces Mixture of Score Guidance (MSG), which decomposes motion into separate components and operates directly on pretrained diffusion models without additional training. Similarly, MotionFlow [Meral2024] utilizes cross-attention maps to facilitate motion transfer during inference, enabling more flexible and generalizable motion synthesis. Furthermore, DiTFlow [Pondaven2024] optimizes Attention Motion Flow (AMF) to transfer motion by utilizing Diffusion Transformers (DiTs) for higher motion consistency. Although these diffusion-based techniques achieve high-fidelity motion synthesis, their dependence on iterative denoising and complex optimization processes results in significant computational overhead, making them less viable for real-time applications.
Crucially, our targeted applications as described in Section 1 require: (i) semantic and spatial consistency tracking the same object across frames while it moves and (ii) geometric consistency preserving stable, physically plausible motion cues. A self-supervised keypoint detector directly supports these requirements [thewlis2017unsupervised, zhang2018unsupervised] by learning to place points at locations which are most informative for predicting how things move, yielding a compact, robust motion signal that is well-suited to forecasting and real-time deployment on resource-constrained platforms. Compared to the other approaches described above, keypoint-based methods like FOMM are more computationally efficient, making them ideal for real-time video applications where low-latency motion transfer is crucial. Our study builds upon keypoint-based motion transfer methods such as the FOMM and demonstrates high generalizability, thereby enabling efficient bandwidth optimization across various keypoint-based architectures such as [siarohin2019animating, siarohin2019first, Wang2021, Zhao2022]. This makes our proposed architecture particularly well-suited for applications such as video reconstruction and animation while ensuring real-time processing capabilities.
@@ 3.1 The Proposed Pipeline
Our real-time motion transfer proposal is based on the FOMM pipeline [siarohin2019first] , although in general our method can be used for realizing bandwidth savings in other keypoint-based architectures such as [Zhao2022, Wang2021]. In the FOMM architecture, inputs to the pipeline consist of a source image $\textbf{S}\in\mathbb{R}^{3\times H\times W}$, where 3 denotes the number of channels corresponding to Red, Green, Blue (RGB), and $H$ and $W$ denote the height and width of the image respectively. and a sequence of $N$ driving video frames $\textbf{D}\in\mathbb{R}^{3\times H\times W}$. At the start of the pipeline there is a keypoint detector which is a convolutional neural network that is used to produce $K$ feature maps from the source image S and each frame of the driving video D. These extracted feature maps are then normalized and condensed to form keypoints by computing the spatial expectation of the maps [Minderer2019]. Each keypoint consists of the coordinates $x,y$ along with parameters of the local affine transformations around each keypoint. The latter set of features model the motion around a keypoint as a 22x22 Jacobian matrix. In this work we set the hyperparameter $K=10$ which gives sufficiently good results. The 20 components representing the coordinates and the 40 components of the Jacobian matrix form a 60 dimensional time series. The keypoint predictor takes $N$ such keypoints corresponding to the driving video frames $\bf D$ and forecasts the next $M$ values. Following this the predicted keypoints go through the dense motion network which is used to align the keypoints captured from source image S and driving video D. The dense motion network is another convolutional neural network that generates heatmaps from the keypoints and then computes the dense motion field. The motion field is basically a function $\hat{\mathcal{T}}_{S\leftarrow D}$ which assigns every pixel position in D to its equivalent position in S. In addition, an occlusion mask $\hat{\mathcal{O}}_{S\leftarrow D}$ is generated by the dense motion network that is responsible for determining which image parts of D must be reconstituted by warping from S and which sections should be infilled based on the context. Finally, the generation module synthesizes the source image S moving in accordance with the driving video D. To achieve this, we utilize a generator network that warps the source image based on $\hat{\mathcal{T}}_{S\leftarrow D}$ and inpaints the occluded regions that are not visible in the source image [siarohin2019first]. In this paper we demonstrate the use of the real-time motion transfer pipeline for multi-step ahead single video frame prediction as well as generating diverse samples of future video frames. The proposed pipeline for single and diverse sample prediction are demonstrated in Figures 1 and 2.
@@ 3.2 Keypoint Prediction Models
GRU was proposed in [cho2014gru] and is a variant of the Recurrent Neural Network (RNN) which is designed to perform forecasting of high dimensional data. The problem of keypoint prediction can also be addressed using Deep Generative Models such as Variational Autoencoder (VAE) [kingma2013auto] and Normalizing Flow (NF) [Rezende2015]. Both of these are generative models that learn a probabilistic representation of high-dimensional data. Although VAEs and NFs are not inherently architected to perform forecasting, in this work we adapt them for keypoint prediction by reconstructing a future value from a given input. Our primary focus in this work is to demonstrate the forecasting capabilities of two generative time series models namely VRNN and GRU-NF. The VRNN model proposed by [Chung2015] combines the strengths of VAEs and RNNs to model sequential data whereas the GRU-NF proposed by [Rasul2020] combines a GRU with an NF. In the following subsections we review the mathematical setup for each of these five models and also discuss how they are trained in our real-time motion transfer pipeline for keypoint forecasting.
@@ 3.2.1 Gated Recurrent Unit (GRU)
Let an input sequence of keypoints be denoted as $\boldsymbol{x}=(\boldsymbol{x}_{1},\boldsymbol{x}_{2},...,\boldsymbol{x}_{t})$. At each timestep $t$, the GRU maintains a hidden state $\boldsymbol{h}_{t}$ that is updated based on the current input $\boldsymbol{x}_{t}$ and the previous hidden state $\boldsymbol{h}_{t-1}$. This update mechanism of the hidden states in GRUs incorporates gating units—specifically, an update gate $\boldsymbol{z}_{t}$ and a reset gate $\boldsymbol{r}_{t}$—which control the flow of information through the network [cho2014gru]. First, the reset gate is computed as:
$\boldsymbol{r}_{t}=\sigma\left(W_{r}\boldsymbol{x}_{t}+U_{r}\boldsymbol{h}_{t-1}\right)$ (1)
where $\sigma(\cdot)$ is the sigmoid function, and $W_{r}$, $U_{r}$ are learnable weight matrices. Using the reset gate, a candidate hidden state $\tilde{\boldsymbol{h}}_{t}$ is then calculated as below:
$\tilde{\boldsymbol{h}}_{t}=\tanh\left(W\boldsymbol{x}_{t}+U\left(\boldsymbol{r}_{t}\odot\boldsymbol{h}_{t-1}\right)\right)$ (2)
@@
$\boldsymbol{z}_{t}=\sigma\left(W_{z}\boldsymbol{x}_{t}+U_{z}\boldsymbol{h}_{t-1}\right)$ (3)
where $W_{z}$, $U_{z}$ are learnable weight matrices. Finally, the hidden state is estimated using the update gate $\boldsymbol{z}_{t}$ and the candidate hidden state $\boldsymbol{\tilde{h}}_{t}$ as follows:
$\boldsymbol{h}_{t}=\left(1-\boldsymbol{z}_{t}\right)\odot\boldsymbol{h}_{t-1}+\boldsymbol{z}_{t}\odot\tilde{\boldsymbol{h}}_{t}$ (4)
This gating mechanism allows the GRU to selectively retain or update information from previous input. During prediction, the GRU maps its hidden states to a probability distribution $p$ of outputs $\boldsymbol{x}_{t+1}$ as below:
$p\left(\boldsymbol{x}_{t+1}\mid\boldsymbol{x}_{\leq t}\right)=g_{\tau}\left(\boldsymbol{h}_{t}\right)$ (5)
where $g_{\tau}$ is a nonlinear function with learnable parameters $\tau$ that transforms the hidden state $\boldsymbol{h}_{t}$ to the corresponding output $\boldsymbol{x}_{t+1}$. In our work, the GRU is trained to predict $\boldsymbol{x}_{t+1}$ from $\boldsymbol{x}_{t}$ by minimizing the mean squared error (MSE) loss for $N$ training samples between the predicted output $\boldsymbol{\hat{x}}$ and its corresponding ground truth $\boldsymbol{x}$ at time $t+1$ as follows:
$\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}(\hat{\boldsymbol{x}}_{t+1}^{(i)}-\boldsymbol{x}_{t+1}^{(i)})^{2}$ (6)
@@ 3.2.2 Variational Autoencoder (VAE)
The VAE is a deep generative model which is trained to minimize the reconstruction error between its inputs and outputs while simultaneously learning a structured, continuous latent space that enables smooth interpolation and sampling of new data. A VAE consists of three main components: an encoder that maps the input data to a distribution over latent variables, a latent space that captures uncertainty through probabilistic representations, and a decoder that reconstructs data using samples drawn from this latent distribution. For a time series of keypoints, a VAE encodes the input $\boldsymbol{x}_{t}$ at each timestep $t$ into a latent representation $\boldsymbol{z}_{t}$ through a probabilistic encoder $q_{\phi}$. The encoder estimates the posterior distribution of the latent variable $\boldsymbol{z}_{t}$ as below:
$q_{\phi}(\boldsymbol{z}_{t}|\boldsymbol{x}_{t})=\mathcal{N}\left(\mu_{\phi}(\boldsymbol{x}_{t}),\sigma_{\phi}^{2}(\boldsymbol{x}_{t})\right)$ (7)
where $\mu_{\phi}(\boldsymbol{x}_{t})$ and $\sigma_{\phi}^{2}(\boldsymbol{x}_{t})$ are the learned mean and variance of the Gaussian distribution. To enable backpropagation during training, the latent variable $\boldsymbol{z}_{t}$ is sampled using the reparameterization trick [rezende2014stochastic]:
$\boldsymbol{z}_{t}=\mu_{\phi}(\boldsymbol{x}_{t})+\sigma_{\phi}(\boldsymbol{x}_{t})\odot\epsilon,\quad\epsilon\sim\mathcal{N}(0,I)$ (8)
where $\epsilon$ is a noise vector drawn from a standard normal distribution and $\odot$ denotes element-wise multiplication. The decoder $p_{\theta}$, parameterized by weights $\theta$, reconstructs the input $\boldsymbol{x}_{t}$ from the latent variable $\boldsymbol{z}_{t}$ by estimating the likelihood as below:
$p_{\theta}(\boldsymbol{x}_{t}|\boldsymbol{z}_{t})=\mathcal{N}\left(\mu_{\theta}(\boldsymbol{z}_{t}),\sigma_{\theta}^{2}(\boldsymbol{z}_{t})\right)$ (9)
Here $p_{\theta}(\boldsymbol{x}_{t}|\boldsymbol{z}_{t})$ represents the likelihood of generating outputs ${\bf x}_{t}$ given the latent variable ${\bf z}_{t}$, the decoder’s output is given by $\mu_{\theta}(\boldsymbol{z}_{t})$ and $\sigma_{\theta}^{2}(\boldsymbol{z}_{t})$ is the variance associated with the reconstruction. Figure 3 illustrates the architecture of a VAE.
@@
In our work, given $k>1$ as the prediction horizon, the VAE is trained to predict $\boldsymbol{x}_{t+k}$ from $\boldsymbol{x}_{t}$ at time $t$ using the loss function as given below:
$\mathcal{L}_{\text{VAE}}=\mathcal{L}_{\text{MSE}}+KL[q_{\phi}(\boldsymbol{z}_{t}|\boldsymbol{x}_{t})||p(\boldsymbol{z}_{t})]$ (10)
@@
$\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}(\hat{\boldsymbol{x}}_{t+k}^{(i)}-\boldsymbol{x}_{t+k}^{(i)})^{2}$ (11)
$KL(q_{\phi}(\bf{z}_{t}|\bf{x}_{t})\,\|\,p(\bf{z}_{t}))$ $=-\frac{1}{2N}\sum_{i=1}^{N}\sum_{j=1}^{J}\Big[1+\log\left(\sigma_{\phi,j}^{2}(\boldsymbol{x}_{t,i})\right)-\mu_{\phi,j}^{2}(\boldsymbol{x}_{t,i})-\sigma_{\phi,j}^{2}(\boldsymbol{x}_{t,i})\Big]$ (12)
The first term in the loss function minimizes the reconstruction loss over $N$ training samples between the predicted output $\boldsymbol{\hat{x}}$ and its corresponding ground truth $\boldsymbol{x}$ at time $t+k$ while the second term, $KL[\cdot||\cdot]$, is the Kullback-Leibler divergence between the approximate posterior $q_{\phi}(\boldsymbol{z}_{t}\mid\boldsymbol{x}_{t})$ and the prior $p(\boldsymbol{z}_{t})$ where $\mathcal{N}(0,I)$ is typically chosen as the prior distribution over the latent space.
During inference, the VAE takes as input a sequence of values from timesteps $t_{1},\dots,t_{n}$, and generates the next $k$ future values by reconstructing each $\boldsymbol{x}_{k+i}$ from $\boldsymbol{x}_{i}$, for $i=1,\dots,n$, where $k>=n$, thereby shifting the input index forward with each prediction.